3x-9 = 4x – 3 3x-9 = 4x – 3 3x-9 = 4x – 3 3x-9 = 4x – 3 3x-9 = 4x – 3 3x-9 = 4x – 3 3x-9 = 4x – 3 3x-9 = 4x – 3

3x-9 = 4x – 3
3x-9 = 4x – 3
3x-9 = 4x – 3
3x-9 = 4x – 3
3x-9 = 4x – 3
3x-9 = 4x – 3
3x-9 = 4x – 3
3x-9 = 4x – 3

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

3x-9 = 4x – 3

3x – 4x = -3 + 9

-x = 6

x = – 6

Thanks